# MobileHolo: A Lightweight Complex-Valued Deformable CNN for High-Quality Computer-Generated Hologram

### v1

@@ V Ablation study
Table IV highlights how performance varies with different initial channel numbers and kernel sizes. The model achieves its highest PSNR of 33.71 dB and SSIM of 0.925 when configured with 10 channels and a kernel size of 3. However, this setup also leads to an increase in the number of parameters and longer inference times. Furthermore, when varying the kernel sizes for deformable convolution while keeping the number of channels fixed at 8, the best results are obtained with a kernel size of 3, whereas the poorest performance is observed with a kernel size of 5.
VI Optical experiment
The results of the optical experiment are presented in Fig. 9. It is clear that the Holo-encoder performs significantly worse than the other models, as it fails to reconstruct detailed information effectively. While HoloNet offers more details compared to the Holo-encoder, it introduces blurring, leading to less clear images. Among the three comparison models, CCNN-CGH shows the highest quality, but suffers from stray light and noise issues. In contrast, our model delivers more consistent reconstruction quality than CCNN-CGH, especially in terms of preserving details.
VII Conclusion
@@ Acknowledgments
Thank all reviewers for their constructive comments.

### v2

@@ 6 Ablation study
Table 4 highlights how performance varies with different initial channel numbers and kernel sizes. The model achieves its highest PSNR of 33.71 dB and SSIM of 0.925 when configured with 10 channels and a kernel size of 3. However, this setup also leads to an increase in the number of parameters and longer inference times. Furthermore, when varying the kernel sizes for deformable convolution while keeping the number of channels fixed at 8, the best results are obtained with a kernel size of 3, whereas the poorest performance is observed with a kernel size of 5.
7 Conclusion
@@
CRediT authorship contribution statement
Shuyang Xie: Writing – original draft, Investigation, Software, Methodology, Validation, Formal analysis, Data curation, Conceptualization. Jie Zhou: Writing – review & editing, Methodology, Validation, Formal analysis, Visualization. Bo Xu: Writing – review & editing, Methodology. Jun Wang: Writing – review & editing, Resource, Supervision. Renjing Xu: Writing – review & editing, Supervision, Resource, Funding acquisition.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Acknowledgments
This work is supported by the Guangzhou Municipal Science and Technology Project (No. 2023A03J0013).
Data Availability
Data will be made available on request.